### v2

# Semi-Blind Strategies for MMSE Channel Estimation Utilizing Generative Priors

Abstract
This paper investigates semi-blind channel estimation for massive multiple-input multiple-output (MIMO) systems. To this end, we first estimate a subspace based on all received symbols (pilot and payload) to provide additional information for subsequent channel estimation. This additional information enhances minimum mean square error (MMSE) channel estimation. Two variants of the linear MMSE (LMMSE) estimator are formulated, where the first one solves the estimation within the subspace, and the second one uses a subspace projection as a preprocessing step. Theoretical derivations show the latter method’s superior estimation performance in terms of mean square error for uncorrelated Rayleigh fading. Further, we provide asymptotical insights on how the proposed MMSE-based channel estimation strategy outperforms the unbiased Cramer-Rao bound. Subsequently, we introduce parameterizations of these semi-blind LMMSE estimators based on two different conditional Gaussian latent models, i.e., the Gaussian mixture model and the variational autoencoder. Both models learn the propagation environment’s underlying channel distribution based on training data and serve as generative priors for our semi-blind channel estimation. Extensive simulations for real-world measurement data and spatial channel models show the proposed methods’ superior performance compared to state-of-the-art semi-blind channel estimators in terms of MSE.
I Introduction
Accurate channel state information (CSI) is crucial for achieving the expected high data rates promised by multiple-input-multiple-output (MIMO) systems [2, 3, 4]. The CSI describes the communication link between transmitter and receiver, characterized by its time-varying and frequency-selective nature, which is prone to rapid changes, making the task of channel estimation complex [5]. As accurate channel estimates are essential for successfully transmitting data, it is at the center of several research efforts [6, 7].
The most widely adopted methods in wireless communications utilize known training or pilot symbols transmitted across the channel using some of the radio resource blocks [8]. Afterward, the receiver uses the observed signals to determine a reliable CSI estimate. As the number of pilots scales with the number of users, the spectral efficiency decreases for a higher number of users as fewer symbols are available for transmitting data. To enhance channel estimation without increasing the number of pilot symbols, various methods have been developed that leverage the information embedded in the observed data symbols at the receiver to infer channel characteristics [9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25]. These methods exploit structure and redundancy within the transmitted data and yield more accurate CSI estimates.
The benefit of semi-blind channel estimation was first studied in [9], where Cramer-Rao bounds (CRBs) for blind, semi-blind, and training-based channel estimation were investigated in the context of single-input-multiple-output (SIMO) systems. In [10, 11], semi-blind channel estimation schemes were introduced based on maximum likelihood (ML) estimation. The respective estimators’ asymptotic performances were also studied in [10] for infinitely long data sequences. Another asymptotic behavior, where the number of antennas grows to infinity, was studied in [12]. Here, the authors identified two interference components in semi-blind channel estimation, which do not vanish even for large numbers of antennas. Early work on improving the least squares (LS) estimator using a semi-blind algorithm was conducted in [13]. Furthermore, in [14], semi-blind and blind channel estimation was studied to enhance the maximum a-posteriori (MAP) channel estimates in massive MIMO systems. These findings are based on favorable propagation, which only holds for large antenna arrays deployed at the base station (BS). In [15], two semi-blind channel estimators are studied based on the expectation maximization (EM) algorithm. The assumption of a Gaussian distribution for the data was verified, leading to a closed-form solution for the E-step. Another iterative framework optimizing the likelihood based on message passing (MP) is used in [16, 17] and references therein. In [18], a data-aided iterative scheme is proposed for orthogonal time frequency space (OTFS) systems by employing affine-precoded superimposed pilots. The performance improvement achieved with these iterative approaches generally requires computationally costly updates. A low complexity iterative LS channel estimation algorithm is proposed in [19] for a massive MIMO turbo-receiver. In [20], the concept of semi-blind channel estimation was adapted for time domain synchronous-orthogonal frequency division multiplexing (OFDM) systems, where in addition to the pseudo noise sequence in the guard interval, the sent OFDM data symbols are exploited for the channel estimation. In [21, 22, 23], different decision-directed frameworks were proposed, which treat reliably decoded data symbols as additional pilots. All of these methods utilize iterative approaches, thus exhibiting large complexities. In [24], peak-power carriers in an OFDM system are selected to eliminate the need to determine reliable data symbols at the receiver. Recently, a diffusion model-based approach for joint channel estimation and detection was proposed in [25], where a diffusion process is constructed that models the joint distribution of the channels and symbols given noisy observations.
Many state-of-the-art estimators fall in the class of unbiased estimators, which are potentially MSE-suboptimal. Instead, in this article, we focus on general, potentially biased, pilot-based estimators that minimize the MSE by utilizing prior information about the channel distribution, and investigate how these estimators can be extended for the semi-blind case. Contrarily to previous works, e.g., [21, 22, 23], we do not focus on using decoded data symbols as additional pilot but instead want to improve the estimators themselves. Although both approaches can be combined. Notably, in [26] a subspace formulation of the MMSE estimator was already used to mitigate pilot contamination in massive MIMO systems. The MMSE estimator is known to be the conditional mean estimator (CME) [27, Ch. 10], which, in general, is intractable and can not be computed in closed form. Recently, powerful approximations based on machine learning were presented in [28, 29, 30, 31, 32]. The benefit of machine learning is to enhance the task at hand by using prior information captured during the learning stage. For a given BS cell environment, the probability density function (PDF) representing potential user channels can be considered valuable prior information. Since this true underlying distribution is unknown, machine learning methods rely on a representative data set, which is assumed to be available at the BS. Based on this data set, the user channels’ PDF can be learned. The first proposal of using a Gaussian mixture model (GMM) to formulate an estimator was done in [28] for the case of image processing. The approaches in [30, 31, 32] build on that by constructing a conditionally Gaussian latent model (CGLM) for the PDF of a BS cell environment. The learned CGLM not only enables MMSE channel estimation in [30, 31, 32] but can also be used for, e.g., a limited feedback scheme as in [33]. In this work, we propose to utilize CGLMs to parameterize the CME in the semi-blind setting.
This work’s contributions are summarized as follows:
  • We introduce two variants of the linear MMSE (MMSE) estimator incorporating subspace knowledge provided by the payload data symbols. Firstly, we depict how the LMMSE channel estimator can solve a subspace estimation problem [26]. As an alternative, we propose a new projection method that is computationally more efficient since it allows for the pre-calculation of LMMSE filters.
  • With theoretical derivations, we show the proposed projection method’s superior MSE performance in the case of uncorrelated Rayleigh fading and perfect subspace knowledge.
  • We provide an asymptotical analysis of how the projected LMMSE outperforms the unbiased CRB for perfect subspaces. These derivations highlight the MMSE-based channel estimation’s superiority compared to unbiased channel estimation in the case of semi-blind channel estimation.
  • We show how the GMM [30] and variational autoencoder (VAE) [32], instances of the class of CGLMs, can be used to parameterize the semi-blind LMMSE estimator.
  • Extensive simulations on different datasets, consisting of typical massive MIMO systems with multiple users and real-world measurement data, show our proposed methods’ superior performance compared to state-of-the-art semi-blind channel estimation algorithms in terms of the MSE.
Preliminary results were presented in [1] and extended to the multi-user MIMO case in [34], which we extend further in the following aspects. The theoretical analyses in Section III enhance the proposed semi-blind channel estimation strategies’ foundation and provide analytic insights into the proposed projection method’s superior performance. We extend our concept of semi-blind MMSE channel estimation to the whole class of CGLMs, providing a more general framework to parameterize the semi-blind LMMSE estimator. Finally, we provide more comprehensive simulation results to show our proposed strategies’ strengths.
Notations: Matrices and vectors are denoted with boldface symbols. $\bm{0}$ and $\mathbf{I}_{N}$ denote the zero vector of appropriate size and the identity matrix of size $N\times N$, respectively. $\mathbb{E}[\cdot]$, $\mathrm{tr}(\cdot)$, and $\mathrm{range}(\cdot)$ denote the expectation, trace, and range operators, respectively. We use $(\cdot)^{\mathrm{T}}$, $(\cdot)^{\mathrm{H}}$, $(\cdot)^{-1}$, $(\cdot)^{\dagger}$ to denote the transpose, conjugate transpose, inverse, and pseudo inverse. $\|\cdot\|$ denotes the $2$-norm of a vector. $\mathcal{N}_{\mathbb{C}}(\bm{\mu},{\bm{C}})$ denotes the circularly symmetric complex Gaussian distribution with mean $\bm{\mu}$ and covariance matrix $\bm{C}$}.
II System and Channel Model
We consider a multi-user system with $J$ single-antenna users and a BS equipped with $M$ receive antennas. We focus on the uplink scenario for this work, but the proposed methods can also be extended to the downlink. The received signal vector at time instance $n$ is then
${\bm{y}}(n)$ $={\bm{H}}{\bm{x}}(n)+{\bm{n}}(n),\quad n=1,...,N,$ (1)
where ${\bm{x}}(n)=[x_{1}(n),...,x_{J}(n)]^{\mathrm{T}}\in\mathbb{C}^{J}$ and ${\bm{n}}(n)\in\mathbb{C}^{M}$ denote the signal sent by each of the $J$ users and the additive noise, respectively, whereas ${\bm{H}}=[{\bm{h}}_{1},...,{\bm{h}}_{J}]$ contains the individual user channels ${\bm{h}}_{j}\in\mathbb{C}^{M}$. The case of multiple antennas at the users can be transformed into (1) by viewing each active stream as a different user with its corresponding channel ${\bm{h}}_{j}$ being the respective effective channel. For further details on semi-blind channel estimation in multi-user MIMO, we refer the reader to [34]. For the case of downlink semi-blind channel estimation, a subspace of the effective channel needs to be identifiable to improve estimation performance. This is the case if not all degrees of freedom are utilized, e.g., fewer pilots than receive antennas. We consider a channel coherence interval larger than the number of snapshots $N$, i.e., the channels are constant over all snapshots. We assume that the noise is Gaussian with ${\bm{n}}(n)\sim\mathcal{N}_{\mathbb{C}}(\bm{0},{\bm{C}}_{\bm{n}}=\sigma^{2}\mathbf{I}_{M})$ and the sent signals satisfy $\mathbb{E}[{\bm{x}}(n){\bm{x}}^{\mathrm{H}}(n)]=\frac{1}{J}\mathbf{I}_{J}$.
In conventional channel estimation schemes, each user’s signals include $N_{p}$ uplink pilots. These pilots are known to the BS. Hence, the received observations at the BS side are
${\bm{Y}}=\left[{\bm{Y}}_{p}^{\prime},{\bm{Y}}_{d}\right]={\bm{H}}\left[{\bm{P}},{\bm{D}}\right]+{\bm{N}}={\bm{H}}{\bm{X}}+{\bm{N}},$ (2)
where ${\bm{Y}}\in\mathbb{C}^{M\times N}$, ${\bm{Y}}^{\prime}_{p}\in\mathbb{C}^{M\times N_{p}}$, ${\bm{Y}}_{d}\in\mathbb{C}^{M\times N-N_{p}}$, ${\bm{P}}\in\mathbb{C}^{J\times N_{p}}$, and ${\bm{D}}=[{\bm{d}}_{1},\dots,{\bm{d}}_{J}]^{\mathrm{T}}$ denote all received observations, received pilot observations, received payload data observations, sent pilots, and sent payload data symbols, respectively. The $j$-th user’s payload symbols ${\bm{d}}_{j}\in\mathbb{S}_{j}^{N-N_{p}}$ include realizations of the used transmit constellation, where $\mathbb{S}_{j}$ denotes the set of constellation points of user $j$. In general, the employed user constellations can differ from each other. In order to fully illuminate the channels, the number of pilots is, at minimum, the number of users $N_{p}\geq J$, and orthogonal pilots are used. We set ${\bm{P}}{\bm{P}}^{\mathrm{H}}=\frac{N_{p}}{J}\mathbf{I}_{M}$, and utilize submatrices of the discrete Fourier transform (DFT) matrix of size $N_{p}\times N_{p}$. After decorrelating the orthogonal pilot sequences, the received pilot observations simplify to
${\bm{Y}}_{p}={\bm{Y}}^{\prime}_{p}{\bm{P}}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\dagger}}={\bm{H}}{\bm{P}}{\bm{P}}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\dagger}}+{\bm{N}}{\bm{P}}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\dagger}}={\bm{H}}+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\sqrt{\frac{J}{N_{p}}}}{\bm{N}}_{p},$ (3)
where ${\bm{N}}_{p}$ has the same statistics as $\bm{N}$} and, hence, we can omit the subscript. This decorrelation lets us consider channel estimation from a per-user perspective in the subsequent discussions. For reasons of simpler readability, the index for the respective user is, therefore, no longer given in the following. Consequently, we denote a user’s pilot observation as
${\bm{y}}_{p}={\bm{h}}+{\bm{n}},$ (4)
with ${\bm{n}}\sim\mathcal{N}_{\mathbb{C}}(\bm{0},{\bm{C}}_{\bm{n}}=\sigma^{2}_{\mathrm{eff}}\mathbf{I}_{M})$, where $\sigma^{2}_{\mathrm{eff}}=\sigma^{2}\frac{J}{N_{p}}$ denotes the effective noise variance. Further, we assume normalized channels as $\mathbb{E}[\|{\bm{h}}\|^{2}]=M$, which lets us define the signal-to-noise ratio (SNR) as $\sigma^{-2}$. As the number of pilots $N_{p}$ results in a simple scaling of SNR in the pilot transmission phase, we use $N_{p}=J$ in the rest of this work.
II-A Spatial Channel Model
We consider a spatial channel model based on [35], where the channel vectors are considered as conditionally Gaussian distributed [29]
${\bm{h}}\mid{\bm{\delta}}\sim\mathcal{N}_{\mathbb{C}}\left(\bm{0},{\bm{C}}_{\bm{\delta}}\right),$ (5)
based on a set of parameters $\bm{\delta}$}, which describe the directions and properties of the multi-path propagation clusters. The central angles are drawn independently from a uniform distribution in $0,2\pi]$, and the path gains are independent zero-mean Gaussians. The spatial covariance matrix is given by
${\bm{C}}_{\bm{\delta}}=\int_{-\pi}^{\pi}g(\vartheta,{\bm{\delta}}){\bm{a}}(\vartheta){\bm{a}}^{\mathrm{H}}(\vartheta)\mathrm{d}\vartheta,$ (6)
where $g(\vartheta,{\bm{\delta}})$ is the power density consisting of a weighted sum of Laplace densities, which have standard deviations $\sigma_{\text{AS}}$ corresponding to the propagation clusters’ angular spread. The BS employs a uniform linear array (ULA) with $M$ antennas and $\lambda/2$ spacing. The steering vector is then given as
${\bm{a}}(\vartheta)=\frac{1}{\sqrt{M}}\left[1,\mathrm{e}^{-\mathrm{j}\pi\sin(\vartheta)},...,\mathrm{e}^{-\mathrm{j}\pi(M-1)\sin(\vartheta)}\right]^{\mathrm{T}}.$ (7)
We consider a new $\bm{\delta}$} for every channel sample and draw the sample according to ${\bm{h}}\sim\mathcal{N}_{\mathbb{C}}\left(\bm{0},{\bm{C}}_{\bm{\delta}}\right)$.
II-B Measurement Campaign
Since synthetic data captures real-world CSI characteristics only up to some extent, we utilize real-world data from a measurement campaign conducted at the Nokia campus in Stuttgart, Germany, in October/November 2017, cf. [36]. The BS antenna with a uniform rectangular array (URA) comprises $N_{v}=4$ vertical ($\lambda$ spacing) and $N_{h}=16$ horizontal ($\lambda/2$ spacing) single polarized patch antennas. The operating carrier frequency is $2.18$ GHz, and the antenna was mounted on a rooftop approximately $20$ meters above the ground. For further details, we refer the reader to [36].
III Semi-Blind Channel Estimation using Perfect Statistical Knowledge
In this section, we introduce two variants of the LMMSE estimator incorporating information provided by the payload data symbols. To do so, we restrict ourselves for now to the case where perfect statistical knowledge is available at the receiver.
In channel estimation, commonly, only the pilot observation ${\bm{y}}_{p}$ is considered for channel estimation. The MSE-optimal estimator given the pilot observation ${\bm{y}}_{p}$ is the CME
$\hat{{\bm{h}}}_{\text{CME}}=\mathbb{E}\left[{\bm{h}}\mid{\bm{y}}_{p}\right].$ (8)
If the channel sample $\bm{h}$} is drawn from a Gaussian distribution according to
${\bm{h}}\sim\mathcal{N}_{\mathbb{C}}(\bm{0},{\bm{C}}),$ (9)
and if further this statistic is known at the receiver, the genie-aided CME can be formulated as [27, Ch. 10]
$\hat{{\bm{h}}}_{\text{CME}}={\bm{C}}\left({\bm{C}}+{\bm{C}}_{\bm{n}}\right)^{-1}{\bm{y}}_{p}.$ (10)
This LMMSE estimator achieves the MSE of
$\mathrm{MSE}^{\mathrm{plain}}=\mathrm{tr}\left[{\bm{C}}-{\bm{C}}({\bm{C}}+{\bm{C}}_{\bm{n}})^{-1}{\bm{C}}\right].$ (11)
For ${\bm{C}}_{\bm{n}}=\sigma^{2}\mathbf{I}_{M}$, the MSE can be expressed using the Woodbury identity as
$\mathrm{MSE}^{\mathrm{plain}}=\sum_{i=1}^{M}\frac{{\rho}_{i}\sigma^{2}}{{\rho}_{i}+\sigma^{2}},$ (12)
where $\rho_{i}$ are the eigenvalues of $\bm{C}$}. In contrast to MMSE-based channel estimation, ML-based channel estimation does not utilize knowledge about the channels’ prior distribution, making it a suboptimal estimation problem.
For our considerations concerning semi-blind channel estimation, we assume knowledge about the subspace defined by $\mathrm{range}({\bm{H}})=\mathrm{range}({\bm{V}})$, where we denote with $\bm{V}$} the left singular vectors of $\bm{H}$}, which span the same subspace as the columns of $\bm{H}$}. Generally, we can formulate the (suboptimal) ML estimate of $\bm{H}$} in view of (2) as [15]
$\min_{{\bm{H}},{\bm{X}}}\sum_{n=1}^{J}\left\|{\bm{y}}(n)-{\bm{h}}_{n}\right\|^{2}+\sum_{n=J+1}^{N}\left\|{\bm{y}}(n)-{\bm{H}}{\bm{x}}(n)\right\|^{2},$ (13)
where the first term belongs to the pilot observations, and the second part refers to the observation obtained from the payload symbols. Obtaining a closed-form solution for (13) is known to be hard [37]. In [15], the EM algorithm is introduced to solve this channel estimation problem in terms of maximum likelihood, whereas, in [11], a semi-blind method is derived based on utilizing the subspace $\mathrm{range}({\bm{H}})=\mathrm{range}({\bm{V}})$. Now, given the subspace $\mathrm{range}({\bm{V}})$, we can reformulate (13) as [11]
$\min_{{\bm{S}},{\bm{X}}}\sum_{n=1}^{J}\left\|{\bm{y}}(n)-{\bm{V}}{\bm{s}}_{n}\right\|^{2}+\hskip-1.0pt\sum_{n=J+1}^{N}\hskip-1.0pt\left\|{\bm{y}}(n)-{\bm{V}}{\bm{S}}{\bm{x}}(n)\right\|^{2},$ (14)
with ${\bm{H}}={\bm{V}}{\bm{S}}$ and ${\bm{S}}=[{\bm{s}}_{1},\dots,{\bm{s}}_{J}]\in\mathbb{C}^{J\times J}$. We can rewrite the second term in (14) as
$\sum_{n=J+1}^{N}\hskip-5.0pt\left\|{\bm{y}}(n)-{\bm{V}}{\bm{S}}{\bm{x}}(n)\right\|^{2}$ $=\hskip-5.0pt\sum_{n=J+1}^{N}\hskip-5.0pt\left\|\left(\mathbf{I}_{M}-{\bm{V}}{\bm{V}}^{\mathrm{H}}\right){\bm{y}}(n)\right\|^{2}$
$\quad+\left\|{\bm{V}}^{\mathrm{H}}{\bm{y}}(n)-{\bm{S}}{\bm{x}}(n)\right\|^{2},$ (15)
where the equality comes from the fact that the two terms are orthogonal to each other. The first term in (15) is constant for given $\bm{V}$}. In the case of a continuous constellation for ${\bm{x}}(n)$, the second term vanishes by solving for $\bm{X}$}. Further, in the case of discrete symmetric signal constellations, the second term in (15) can not be jointly solved for $\bm{S}$} and $\bm{X}$} due to phase ambiguity. Thus, the ML problem for the user of interest reduces to the pilot observations as [11]
$\min_{\bm{s}}\|{\bm{y}}_{p}-{\bm{V}}{\bm{s}}\|^{2},$ (16)
with the closed form solution $\hat{{\bm{h}}}_{\text{ML}}={\bm{V}}{\bm{V}}^{\mathrm{H}}{\bm{y}}_{p}$. This estimator’s MSE is
$\mathrm{MSE}^{\mathrm{ML}}=\mathbb{E}\left[\|{\bm{h}}-{\bm{V}}{\bm{V}}^{\mathrm{H}}{\bm{y}}_{p}\|^{2}\right]=J\sigma^{2}.$ (17)
In the following, we introduce two channel estimation strategies combining the information provided by $\bm{C}$} and $\mathrm{range}({\bm{V}})$.
III-A Subspace Channel Estimator
Using the information in $\mathrm{range}({{\bm{V}}})$, we can solve the estimation within the subspace as previously proposed in [26]. For this, the pilot system model in (4) is mapped into the $J$-dimensional subspace as
${\bm{y}}^{\prime}={\bm{V}}^{\mathrm{H}}{\bm{y}}_{p}$ $={\bm{V}}^{\mathrm{H}}{\bm{h}}+{\bm{V}}^{\mathrm{H}}{\bm{n}}={\bm{h}}^{\prime}+{\bm{n}}^{\prime}.$ (18)
Under the assumption that $\bm{V}$} is chosen independently of $\bm{h}$}, the distribution ${\bm{h}}^{\prime}\sim\mathcal{N}_{\mathbb{C}}(\bm{0},{\bm{V}}^{\mathrm{H}}{\bm{C}}{\bm{V}})$ can be used to formulate the estimate [26]
$\hat{{\bm{h}}}^{\prime}=\;$ ${\bm{V}}^{\mathrm{H}}{\bm{C}}{\bm{V}}\left({\bm{V}}^{\mathrm{H}}{\bm{C}}{\bm{V}}+\sigma^{2}\mathbf{I}_{J}\right)^{-1}{\bm{V}}^{\mathrm{H}}{\bm{y}}_{p}.$ (19)
By design, $\bm{V}$} depends on $\bm{h}$} and, hence, (19) is a suboptimal but feasible estimate for $\bm{h}$$^{\prime}$. After solving the estimation in the subspace for $\bm{h}$$^{\prime}$, the solution can be transformed back using [26]
$\hat{{\bm{h}}}_{\text{sub}}={\bm{V}}\hat{{\bm{h}}}^{\prime}.$ (20)
III-B Projected Channel Estimator
As an alternative approach, we propose using the orthogonal subspace projection ${\bm{P}}_{\bm{H}}={\bm{V}}{\bm{V}}^{\mathrm{H}}$ as a preprocessing filter. Since the projector ${\bm{P}}_{\bm{H}}$ does not affect $\bm{h}$}, the resulting projected pilot observation is given by
$\tilde{{\bm{y}}}={\bm{P}}_{\bm{H}}{\bm{y}}_{p}={\bm{h}}+{\bm{P}}_{\bm{H}}{\bm{n}}={\bm{h}}+\tilde{{\bm{n}}}$. (21)
To formulate the CME $\hat{{\bm{h}}}=\mathbb{E}\left[{\bm{h}}\mid\tilde{{\bm{y}}}\right$], we calculate the covariance matrix of $\tilde{{\bm{y}}}$ as
$\mathbb{E}\left[\tilde{{\bm{y}}}\tilde{{\bm{y}}}^{\mathrm{H}}\right]$ $={\bm{C}}+\mathbb{E}\left[\sigma^{2}{\bm{P}}_{\bm{H}}\right],$ (22)
where the mixing terms vanish due to the independence of $\bm{h}$} and $\bm{n}$}. To get an intuitive understanding of the noise covariance in (22), let us consider a scenario involving spatially uncorrelated channels, meaning that path gains and channel directions are uncorrelated. This is the case when users are uniformly distributed over the directions, e.g., the spatial channel model in Section II-A, resulting in a scenario’s channel covariance matrix that is a scaled identity [38, Def. 2.3]. In such a case, the matrices with the sample covariance matrix’s eigenvectors are distributed with Haar measure [39, Chap. 1], i.e., uniformly distributed on the manifold of unitary matrices. Assuming spatially uncorrelated channels we have
${\bm{C}}_{\tilde{{\bm{n}}}}=\sigma^{2}\frac{J}{M}\mathbf{I}_{M},$ (23)
which we assume to hold for the remainder of this section. We can then formulate the projected LMMSE estimator as
$\hat{{\bm{h}}}_{\text{proj}}={\bm{C}}\left({\bm{C}}+{\bm{C}}_{\tilde{{\bm{n}}}}\right)^{-1}\tilde{{\bm{y}}}$. (24)
III-C Performance Analysis for Perfect Subspace Knowledge
If (23) is true, the MSE of the proposed projected channel estimator can directly be written as (cf. Section -A)
$\mathrm{MSE}^{\mathrm{proj}}$ $=\mathrm{tr}\left({{\bm{C}}}-{{\bm{C}}}\left({{\bm{C}}}+\sigma^{2}\frac{J}{M}\mathbf{I}_{M}\right)^{-1}{{\bm{C}}}\right)$ (25)
$=\sum_{i=1}^{M}\frac{\rho_{i}\sigma^{2}}{\frac{M}{J}\rho_{i}+\sigma^{2}}.$ (26)
Comparing the performance to the plain LMMSE from (12) we see that
$\frac{\rho_{i}\sigma^{2}}{\frac{M}{J}\rho_{i}+\sigma^{2}}\leq\frac{\rho_{i}\sigma^{2}}{\rho_{i}+\sigma^{2}},$ (27)
holds for every $i\in\{1,\dots,M\}$, resulting in $\mathrm{MSE}^{\mathrm{proj}}\leq\mathrm{MSE}^{\mathrm{plain}}$. The inequality in (27) only holds with equality if $J=M$. Additionally, we can compare the MSE of the projected LMMSE to the ML estimator from (17) by reformulating (26) as
$\mathrm{MSE}^{\mathrm{proj}}=\frac{J}{M}\sigma^{2}\sum_{i=1}^{M}\frac{\rho_{i}}{\rho_{i}+\frac{J}{M}\sigma^{2}}\leq J\sigma^{2}=\mathrm{MSE}^{\text{ML}}.$ (28)
In the case of uncorrelated Rayleigh fading, the channel covariance matrix is given as ${\bm{C}}=\mathbf{I}_{M}$. The projected variant’s MSE results in
$\mathrm{MSE}^{\mathrm{proj}}_{\mathrm{iid}}=\frac{JM\sigma^{2}}{{M}+J\sigma^{2}}.$ (29)
Now let us compare the projected variant to the subspace LMMSE. For the case of ${\bm{C}}=\mathbf{I}_{M}$, the subspace LMMSE estimator boils down to
$\hat{{\bm{h}}}_{\mathrm{sub}}=\frac{1}{1+\sigma^{2}}{\bm{V}}{\bm{V}}^{\mathrm{H}}{\bm{y}},$ (30)
with its corresponding MSE as (cf. Section -B)
$\mathrm{MSE}^{\mathrm{sub}}_{\mathrm{iid}}=\frac{\sigma^{2}(M\sigma^{2}+J)}{(1+\sigma^{2})^{2}}\geq\mathrm{MSE}^{\mathrm{proj}}_{\mathrm{iid}}.$ (31)
Fig. 1a shows the individual channel estimators’ performances based on perfect subspace knowledge for the case of uncorrelated Rayleigh fading. As one can see, the projected channel estimator outperforms all other estimators across the whole SNR range. Further, we realize that the subspace LMMSE converges to the ML method from above.
Additionally, performance guarantees of the proposed projected LMMSE compared to the unbiased CRB as derived in [15] can be given. We formulate the following theorem for the asymptotic region, where $M$ and $N$ tend to infinity.
Theorem 1.
We consider the limits of $N\to\infty$ and $M/N\to\alpha$ with $\alpha>0$, while keeping the number of pilot symbols in the order of users $J$. Then, the projected LMMSE based on perfect knowledge of the subspace spanned by the left singular vectors of the channel matrix $\bm{H}$} is equal to or better than any unbiased estimator in terms of MSE for uncorrelated Rayleigh fading.
Proof.
We take the limit of the MSE of the projected LMMSE estimator as
$\lim_{\begin{subarray}{c}N\to\infty\\ \frac{M}{N}\to\alpha\end{subarray}}\mathrm{MSE}^{\mathrm{proj}}$ $=\lim_{\begin{subarray}{c}N\to\infty\\ \frac{M}{N}\to\alpha\end{subarray}}\frac{J}{M}\sigma^{2}\sum_{i=1}^{M}\frac{\rho_{i}}{\rho_{i}+\frac{J}{M}\sigma^{2}}$ (32)
$=J\sigma^{2}\leq(1+\alpha)J\sigma^{2}=\mathrm{CRB}_{\mathrm{iid,d}},$ (33)
where $\mathrm{CRB}_{\mathrm{iid,d}}$ denotes the deterministic CRB’s limit for uncorrelated Rayleigh fading. The last equality in (33) was proven in [40, Theorem 4]. Furthermore, it is stated in [40] that in this asymptotic limit, the deterministic CRB is always lower than the stochastic CRB. ∎
Theorem 1 provides some insights on the asymptotic superiority of our proposed projected LMMSE strategy based on perfect subspace knowledge compared to any unbiased channel estimator. Additionally, the MSE of the subspace variant as well as the plain LMMSE approach infinity for the considered limit, making them irrelevant for this asymptotic regime. Interestingly, the projected LMMSE also approaches $J\sigma^{2}$ for the case where only $M$ tends to infinity, making the projected LMMSE a valid candidate for large antenna systems as the performance no longer depends on $M$.
III-D Maximum Likelihood Subspace Estimation
After introducing the methods utilizing the additional subspace information provided by $\mathrm{range}({\bm{V}})$ to enhance the CSI estimation quality, let us consider the estimation of such a subspace. As the received data symbols are transmitted over the same channels, we can use these payload symbols to estimate the subspace containing all user channels.
To this end, let us reconsider the ML estimate of $\bm{H}$} in (13). Instead of directly optimizing on this ML formulation as done in [15], which generally does not result in the MMSE, we only take this log-likelihood formulation as an intermediate step to estimate the subspace $\mathrm{range}({\bm{V}})$. First, let us consider the payload ML objective formulation in (15). We can then reformulate the problem, neglecting the second term due to phase ambiguity, resulting in [11]
$\max_{\bm{H}}\;\mathrm{tr}\left({\bm{P}}_{\bm{H}}\hat{{\bm{C}}}^{(d)}_{{\bm{y}}\mid{\bm{H}}}\right)$, (34)
where ${\bm{P}}_{\bm{H}}={\bm{H}}({\bm{H}}^{\mathrm{H}}{\bm{H}})^{-1}{\bm{H}}^{\mathrm{H}}={\bm{V}}{\bm{V}}^{\mathrm{H}}$ and $\hat{{\bm{C}}}^{(d)}_{{\bm{y}}\mid{\bm{H}}}=\frac{1}{N-J}{\bm{Y}}_{d}{\bm{Y}}_{d}^{\mathrm{H}}$, with ${\bm{Y}}_{d}$ from (2). The maximization in (34) is solved by setting ${\bm{P}}_{\bm{H}}$ equal to $\hat{{\bm{V}}}_{d}\hat{{\bm{V}}}_{d}^{\mathrm{H}}$ with $\hat{{\bm{V}}}_{d}$ holding the $J$ dominant eigenvectors of the receive sample covariance matrix $\hat{{\bm{C}}}^{(d)}_{{\bm{y}}\mid{\bm{H}}}$. Additionally, it is trivial to see that the first term in (13) is minimized by ${\bm{h}}_{n}={\bm{y}}(n)$. The subspace spanned by the solution ${\bm{h}}_{n}={\bm{y}}(n)$ is the same as the subspace spanned by the $J$ eigenvectors of the sample covariance matrix $\hat{{\bm{C}}}^{(p)}_{{\bm{y}}\mid{\bm{H}}}=\frac{1}{J}{\bm{Y}}_{p}{\bm{Y}}_{p}^{\mathrm{H}}$, which ignores the additional phase information contained in the pilot observation. Thus, the overall subspace estimate $\hat{{\bm{V}}}=[{\bm{v}}_{1},\dots,{\bm{v}}_{J}]$ is found by taking the $J$ dominant eigenvectors ${\bm{v}}_{j}$ of the sample covariance matrix defined as
$\hat{{\bm{C}}}_{{\bm{y}}\mid{\bm{H}}}=\frac{1}{N}{\bm{Y}}{\bm{Y}}^{\mathrm{H}}.$ (35)
This result has also been used in [26].
To save computational complexity while calculating the sample covariance matrix in (35), information from the previous coherence intervals can be utilized. This can be done by adaptively updating the subspace using efficient tracking algorithms as proposed in, e.g., [41, 42].
(a) Perfect Subspace Knowledge
(b) Subspace Estimation based on $N=200$ Snapshots
Figure 1: MSE over the SNR for given channel estimations based on perfect statistical knowledge in a $J=8$ user and $M=64$ antennas scenario with uncorrelated Rayleigh fading and Gaussian symbols using (a) perfect subspace knowledge and (b) $N=200$ snapshots for subspace estimation.
III-E Performance Analysis for Estimated Subspace
Fig. 1b shows the individual channel estimators’ performances based on an estimated subspace using $N=200$ snapshots for uncorrelated Rayleigh fading. Additionally, the unbiased CRBs derived in [15] and their extensions to the Bayesian Cramer-Rao bounds (BCRBs), derived in Section -C, are shown. For increasing SNR, all semi-blind channel estimators considered in this section converge to the CRBs and BCRBs. Furthermore, we see that for low SNR values the introduced MMSE-based semi-blind channel estimators outperform the different unbiased CRBs, which are lower bounds for, e.g., the EM-based estimator from [15].
In order to quantify the error introduced by the subspace estimation in comparison to perfect subspace knowledge, let us consider the projected LMMSE as introduced in Section III-B in combination with the subspace estimation from Section III-D. We assume Gaussian distributed symbols for the following derivations. The introduced error by estimating the subspace can be bounded using Theorem 2.
Theorem 2.
Let the subspace spanned by the left singular vectors of the channel matrix $\bm{H}$} be estimated based on $N$ observations. The probability that the difference between the projected LMMSE channel estimators utilizing the estimated and true subspace, respectively, is less than $\varepsilon$ is given as
$P\left(\mathbb{E}_{{\bm{X}},{\bm{N}}}[\|\hat{{\bm{h}}}_{\mathrm{proj}}(\hat{{\bm{V}}})-\hat{{\bm{h}}}_{\mathrm{proj}}({{\bm{V}}})\|^{2}]\leq\varepsilon\right)\\ \geq 1-\frac{4k^{2}(M-J)\lambda_{\max}({\bm{W}}_{\mathrm{proj}}^{\mathrm{H}}{\bm{W}}_{\mathrm{proj}})}{J^{2}N\varepsilon}\sum_{j\leq J}\frac{|s_{\max}|^{2}+2\sigma^{2}}{\lambda_{j}^{2}},$ (36)
where $\lambda_{j}$ is the $j$-th eigenvalue of ${\bm{H}}{\bm{H}}^{\mathrm{H}}$, $s_{\max}=\max_{j\in\{1,\dots,J\}}s_{j}$ with ${\bm{h}}={\bm{V}}{\bm{s}}$, and
$k^{2}=\sigma^{2}\left(\sigma^{2}+\mathrm{tr}({\bm{C}}_{{\bm{y}}\mid{\bm{H}}})\right).$ (37)
Proof.
See Appendix -D
The probability bound in Theorem 2 is very loose, as the used inequalities are not tight. Nevertheless, this bound can be used to show the superiority of the projected LMMSE based on an estimated subspace for a high number of observations compared to the other considered estimators. Table I gives the bound values for the case of $N=10^{4}$ observations at SNR =0=0 dB for different number of users. As one can see, the projected LMMSE outperforms the other estimators with high probability.
TABLE I: Lower bound on the probability of the projected LMMSE outperforming the other estimators for $N=10^{4}$ observations at SNR =0=0dB.
Number of users $J$ 88 $16$ $32$
$P(\mathrm{MSE}^{\mathrm{proj}}_{\mathrm{iid}}\leq\mathrm{MSE}^{\mathrm{sub}}_{\mathrm{iid}})$ $\geq 0.9873$ $\geq 0.9791$ $\geq 0.9642$
$P(\mathrm{MSE}^{\mathrm{proj}}_{\mathrm{iid}}\leq\mathrm{MSE}^{\mathrm{ML}}_{\mathrm{iid}})$ $\geq 0.8453$ $\geq 0.9539$ $\geq 0.9907$
$P(\mathrm{MSE}^{\mathrm{proj}}_{\mathrm{iid}}\leq\mathrm{MSE}^{\mathrm{plain}}_{\mathrm{iid}})$ $\geq 0.9946$ $\geq 0.9923$ $\geq 0.9907$
IV Proposed Semi-Blind Channel Estimation - Utilizing Generative Prior
In practice, the user channels do not follow a Gaussian distribution. Hence, (10) can not be utilized directly and the CME given the pilot observation ${\bm{y}}_{p}$ is formulated as
$\hat{{\bm{h}}}_{\text{CME}}=\mathbb{E}\left[{\bm{h}}\mid{\bm{y}}_{p}\right]=\int{\bm{h}}\frac{p_{{\bm{n}}}({\bm{y}}_{p}-{\bm{h}})p({\bm{h}})}{p({\bm{y}}_{p})}\mathrm{d}{\bm{h}}.$ (38)
As seen in (38), the CME generally can not be computed analytically. First, the CME needs access to $p({\bm{h}})$, which is generally unavailable in practice. Additionally, no closed-form solution exists to the integral in (38).
In order to reformulate the CME, we first use the property that for any arbitrarily distributed random variable $\bm{h}$}, we can always find a condition $\bm{c}$} which makes the conditional distribution Gaussian. Secondly, it has been shown in [43] that for wireless communication channels, this conditional Gaussian distribution preserves the zero-mean property as
${\bm{h}}\mid{\bm{\delta}}\sim\mathcal{N}_{\mathbb{C}}(\bm{0},{\bm{C}}_{{\bm{h}}\mid{\bm{c}}}).$ (39)
Thus, we can reformulate the CME as
$\mathbb{E}\left[{\bm{h}}\mid{\bm{y}}_{p}\right]$ $=\mathbb{E}\left[\mathbb{E}\left[{\bm{h}}\mid{\bm{y}}_{p},{\bm{c}}\right]\mid{\bm{y}}_{p}\right]$ (40)
$=\int\mathbb{E}\left[{\bm{h}}\mid{\bm{y}}_{p},{\bm{c}}\right]p({\bm{c}}\mid{\bm{y}}_{p})\mathrm{d}{\bm{c}}$ (41)
$=\int\hat{{\bm{h}}}_{\bm{c}}({\bm{y}}_{p})p({\bm{c}}\mid{\bm{y}}_{p})\mathrm{d}{\bm{c}}$ (42)
where
$\hat{{\bm{h}}}_{\bm{c}}({\bm{y}}_{p})={\bm{C}}_{{\bm{h}}\mid{\bm{c}}}\left({\bm{C}}_{{\bm{h}}\mid{\bm{c}}}+{\bm{C}}_{\bm{n}}\right)^{-1}{\bm{y}}_{p},$ (43)
denotes the LMMSE estimate given $\bm{c}$}. However, finding a suitable condition $\bm{c}$} can be challenging, particularly as the true distribution of $\bm{h}$} is unknown. To this end, CGLMs were proposed in [30, 31, 32], which approximate the CME based on a GMM, mixtures of factor analyzers (MFA), and VAE, respectively. All three methods learn a model that provides the conditional Gaussian distribution ${\bm{h}}\mid{\bm{c}}$ based on a discrete (GMM, MFA) or continuous (VAE) latent variable $\bm{c}$}. This work focuses on the GMM and VAE, which we adapt to semi-blind channel estimation in the following.
IV-A GMM-based Semi-blind Channel Estimation
Algorithm 1 Subspace GMM Channel Estimator
Offline Training Phase
1:Training dataset $\mathcal{H}=\{{\bm{h}}_{t}\}_{t=1}^{T}$
2:Fit the GMM with the EM algorithm, cf. [30]
3:
[…]
8:for $j=1,\dots,J$ do
9:  ${\bm{y}}^{\prime}\leftarrow{\bm{V}}^{\mathrm{H}}{\bm{y}}_{p,j}$
10:  for $k=1,\dots,K$ do
11:   $\hat{{\bm{h}}}^{\prime}_{k}\leftarrow{\bm{V}}^{\mathrm{H}}{\bm{C}}_{k}{\bm{V}}\left({\bm{V}}^{\mathrm{H}}{\bm{C}}_{k}{\bm{V}}+\sigma^{2}\mathbf{I}_{J}\right)^{-1}{\bm{y}}^{\prime}$
12:  end for
13:  $\hat{{\bm{h}}}_{j}\leftarrow{\bm{V}}\sum_{k=1}^{K}p(k\mid{\bm{y}}^{\prime})\hat{{\bm{h}}}^{\prime}_{k}$
14:end for
15:return $\hat{{\bm{h}}}_{j},\forall j=1,\dots,J$
Algorithm 2 Projected GMM Channel Estimator
Offline Training Phase
1:Training dataset $\mathcal{H}=\{{\bm{h}}_{t}\}_{t=1}^{T}$
2:Fit the GMM with the EM algorithm, cf. [30]
3:
[…]
8:for $j=1,\dots,J$ do
9:  $\tilde{{\bm{y}}}\leftarrow{\bm{V}}{\bm{V}}^{\mathrm{H}}{\bm{y}}_{p,j}$
10:  for $k=1,\dots,K$ do
11:   $\hat{{\bm{h}}}_{k}\leftarrow{\bm{C}}_{k}\left({\bm{C}}_{k}+{\bm{C}}_{\tilde{{\bm{n}}}}\right)^{-1}\tilde{{\bm{y}}}$
12:  end for
13:  $\hat{{\bm{h}}}_{j}\leftarrow\sum_{k=1}^{K}p(k\mid\tilde{{\bm{y}}})\hat{{\bm{h}}}_{k}$
14:end for
15:return $\hat{{\bm{h}}}_{j},\forall j=1,\dots,J$
Based on the universal approximation property of GMMs [44], the PDF of $\bm{h}$} is approximated by
$f_{\bm{h}}^{(K)}({\bm{h}})=\sum_{k=1}^{K}p(k)\mathcal{N}_{\mathbb{C}}({\bm{h}};\bm{\mu}_{k},{\bm{C}}_{k}),$ (44)
where $p(k)$, $\bm{\mu}_{k}$, and ${\bm{C}}_{k}$ are the mixing coefficients, means, and covariance matrices of the $k$-th GMM component, respectively. As we consider wireless channels, each component’s mean is set to $\bm{\mu}_{k}=\bm{0}$, cf. [43]. The fitting of the components in (44) is accomplished with the well-known EM algorithm [45] based on a set $\mathcal{H}=\{{\bm{h}}_{t}\}^{T}_{t=1}$ of $T$ channel samples as training data. Based on the formulation in (44) the conditional PDF is
${\bm{h}}\mid k\sim\mathcal{N}_{\mathbb{C}}({\bm{h}};\bm{0},{\bm{C}}_{k}).$ (45)
Thus, we have a discrete latent variable in the case of a GMM, which helps us parameterize the CME. The resulting semi-blind subspace GMM can be formulated as
$\hat{{\bm{h}}}_{\text{sub. GMM}}={\bm{V}}\hat{{\bm{h}}}^{\prime}_{\text{GMM}}={\bm{V}}\sum_{k=1}^{K}p(k\mid{\bm{y}}^{\prime})\hat{{\bm{h}}}^{\prime}_{\text{GMM},k},$ (46)
with
$\hat{{\bm{h}}}^{\prime}_{\text{GMM},k}=\;$ ${\bm{V}}^{\mathrm{H}}{\bm{C}}_{k}{\bm{V}}\left({\bm{V}}^{\mathrm{H}}{\bm{C}}_{k}{\bm{V}}+\sigma^{2}\mathbf{I}_{J}\right)^{-1}{\bm{V}}^{\mathrm{H}}{\bm{y}}_{p},$ (47)
and the corresponding responsibilities
$p(k\mid{\bm{y}}^{\prime})=\frac{p(k)\mathcal{N}_{\mathbb{C}}\left({\bm{y}}^{\prime};\bm{0},{\bm{V}}^{\mathrm{H}}{\bm{C}}_{k}{\bm{V}}+\sigma^{2}\mathbf{I}_{J}\right)}{\sum_{i=1}^{K}p(i)\mathcal{N}_{\mathbb{C}}\left({\bm{y}}^{\prime};\bm{0},{\bm{V}}^{\mathrm{H}}{\bm{C}}_{i}{\bm{V}}+\sigma^{2}\mathbf{I}_{J}\right)}.$ (48)
The projected GMM is
$\hat{{\bm{h}}}_{\text{proj. GMM}}=\sum_{k=1}^{K}p(k\mid\tilde{{\bm{y}}})\hat{{\bm{h}}}_{\text{proj. GMM},k},$ (49)
with
$\hat{{\bm{h}}}_{\text{proj. GMM},k}={\bm{C}}_{k}\left({\bm{C}}_{k}+{\bm{C}}_{\tilde{{\bm{n}}}}\right)^{-1}\tilde{{\bm{y}}}$ (50)
and the associated responsibilities
$p(k\mid\tilde{{\bm{y}}})=\frac{p(k)\mathcal{N}_{\mathbb{C}}\left(\tilde{{\bm{y}}};\bm{0},{\bm{C}}_{k}+{\bm{C}}_{\tilde{{\bm{n}}}}\right)}{\sum_{i=1}^{K}p(i)\mathcal{N}_{\mathbb{C}}\left(\tilde{{\bm{y}}};\bm{0},{\bm{C}}_{i}+{\bm{C}}_{\tilde{{\bm{n}}}}\right)}$. (51)
The respective estimators are summarized in Algorithm 1 and Algorithm 2.
IV-B VAE-based Semi-blind Channel Estimation
Figure 2: Structure of a semi-blind VAE. The matrix $\hat{{\bm{V}}}$ contains the $J$ dominant eigenvectors of (35). The encoder and decoder represent DNNs.
[…]
$\mathcal{L}_{{\bm{\theta}},{\bm{\phi}}}=\mathbb{E}_{q_{\bm{\phi}}}[\log p_{\bm{\theta}}({\bm{h}}\mid{\bm{z}})]-\mathrm{D}_{\mathrm{KL}}(q_{\bm{\phi}}({\bm{z}}\mid\tilde{{\bm{y}}})\mid\mid p({\bm{z}})),$ (52)
[…]
$\mathrm{D}_{\mathrm{KL}}(q_{\bm{\phi}}({\bm{z}}\mid\tilde{{\bm{y}}})\mid\mid p({\bm{z}}))=\mathbb{E}_{q_{\bm{\phi}}}\left[\log\left(\frac{q_{\bm{\phi}}({\bm{z}}\mid\tilde{{\bm{y}}})}{p({\bm{z}})}\right)\right].$ (53)
In the VAE framework, the ELBO is optimized using deep neural networks (DNNs) and the reparameterization trick [46]. In order to do so, the involved distributions are defined as
$p({\bm{z}})$ $=\mathcal{N}(\bm{0},\mathbf{I}_{Z}),$
$p_{\bm{\theta}}({\bm{h}}\mid{\bm{z}})$ $=\mathcal{N}_{\mathbb{C}}(\bm{\mu}_{\bm{\theta}}({\bm{z}}),{{\bm{C}}}_{\bm{\theta}}({\bm{z}})),$ (54)
[…]
The resulting semi-blind VAE structure is shown in Fig. 2. In the case of a ULA or URA at the BS, the channel covariance matrix is either Toeplitz or block-Toeplitz, respectively. As shown in [43], the conditional covariance matrix at the VAE’s output preserves this structure. Thus, we parameterize the output covariance matrix as
${{\bm{C}}}_{\bm{\theta}}({\bm{z}})={\bm{Q}}^{\mathrm{H}}\mathrm{diag}({\bm{c}}_{\bm{\theta}}({\bm{z}})){\bm{Q}},$ (55)
where ${\bm{Q}}={\bm{Q}}_{M}$ or ${\bm{Q}}={\bm{Q}}^{\prime}_{N_{v}}\otimes{\bm{Q}}^{\prime}_{N_{h}}$, respectively, where ${\bm{Q}}_{M}$ is a DFT matrix of size $M$ resulting in a circulant approximation, cf. [32], and ${\bm{Q}}^{\prime}_{N_{x}}$ contains the first $N_{x}$ columns of the $2N_{x}\times 2N_{x}$ DFT matrix resulting in a block-Toeplitz parameterization, cf. [47]. Further, for the (block-)Toeplitz parameterization, we can set $\bm{\mu}_{\bm{\theta}}({\bm{z}})=\bm{0}$, cf. [43].
Algorithm 3 Subspace VAE Channel Estimator
Offline Training Phase
1:Training dataset $\mathcal{H}=\{{\bm{h}}_{t}\}_{t=1}^{T}$
2:Fit the VAE by optimizing the ELBO, cf. [32]
3:
[…]
8:for $j=1,\dots,J$ do
9:  $\tilde{{\bm{y}}}\leftarrow{\bm{V}}{\bm{V}}^{\mathrm{H}}{\bm{y}}_{p,j}$
[…]
11:  $\hat{{\bm{h}}}_{j}\leftarrow{\bm{V}}{\bm{V}}^{\mathrm{H}}{\bm{C}}_{\bm{\theta}}({\bm{z}}){\bm{V}}\left({\bm{V}}^{\mathrm{H}}{\bm{C}}_{\bm{\theta}}({\bm{z}}){\bm{V}}+\sigma^{2}\mathbf{I}_{J}\right)^{-1}$
12:     $\times({\bm{V}}^{\mathrm{H}}{\bm{y}}_{p}-{\bm{V}}^{\mathrm{H}}\bm{\mu}_{\bm{\theta}}({\bm{z}}))-{\bm{V}}{\bm{V}}^{\mathrm{H}}\bm{\mu}_{\bm{\theta}}({\bm{z}})$
13:end for
14:return $\hat{{\bm{h}}}_{j},\forall j=1,\dots,J$
Algorithm 4 Projected VAE Channel Estimator
Offline Training Phase
1:Training dataset $\mathcal{H}=\{{\bm{h}}_{t}\}_{t=1}^{T}$
2:Fit the VAE by optimizing the ELBO, cf. [32]
3:
[…]
8:for $j=1,\dots,J$ do
9:  $\tilde{{\bm{y}}}\leftarrow{\bm{V}}{\bm{V}}^{\mathrm{H}}{\bm{y}}_{p,j}$
[…]
11:  $\hat{{\bm{h}}}_{j}\leftarrow\bm{\mu}_{\bm{\theta}}({\bm{z}})+{\bm{C}}_{\bm{\theta}}({\bm{z}})\left({\bm{C}}_{\bm{\theta}}({\bm{z}})+{\bm{C}}_{\tilde{{\bm{n}}}}\right)^{-1}(\tilde{{\bm{y}}}-\bm{\mu}_{\bm{\theta}}({\bm{z}}))$
12:     $\times(\tilde{{\bm{y}}}-\bm{\mu}_{\bm{\theta}}({\bm{z}}))$
13:end for
14:return $\hat{{\bm{h}}}_{j},\forall j=1,\dots,J$
After successfully training the VAE, the output is a local parameterization of $p({\bm{h}})$ as conditionally Gaussian
${\bm{h}}\mid{\bm{z}}\sim p_{\bm{\theta}}({\bm{h}}\mid{\bm{z}}).$ (56)
As analyzed in [32], it is a reasonable approximation to set
$p({\bm{z}}\mid\tilde{{\bm{y}}})=\begin{cases}1\quad\text{if }{\bm{z}}=\bm{\mu}_{\bm{\phi}}(\tilde{{\bm{y}}}),\\ 0\quad\text{otherwise}.\end{cases}$ (57)
Based on this parameterization, we can formulate the semi-blind VAE-based estimators as
$\hat{{\bm{h}}}_{\text{proj. VAE}}=\;$ $\bm{\mu}_{\bm{\theta}}({\bm{z}})+{\bm{C}}_{\bm{\theta}}({\bm{z}})\left({\bm{C}}_{\bm{\theta}}({\bm{z}})+{\bm{C}}_{\tilde{{\bm{n}}}}\right)^{-1}(\tilde{{\bm{y}}}-\bm{\mu}_{\bm{\theta}}({\bm{z}})),$ (58)
and
$\hat{{\bm{h}}}_{\text{sub. VAE}}=\;$ ${\bm{V}}{\bm{V}}^{\mathrm{H}}{\bm{C}}_{\bm{\theta}}({\bm{z}}){\bm{V}}\left({\bm{V}}^{\mathrm{H}}{\bm{C}}_{\bm{\theta}}({\bm{z}}){\bm{V}}+\sigma^{2}\mathbf{I}_{J}\right)^{-1}$
$\times({\bm{V}}^{\mathrm{H}}{\bm{y}}_{p}-{\bm{V}}^{\mathrm{H}}\bm{\mu}_{\bm{\theta}}({\bm{z}}))-{\bm{V}}{\bm{V}}^{\mathrm{H}}\bm{\mu}_{\bm{\theta}}({\bm{z}}).$ (59)
The respective estimators are summarized in Algorithm 3 and Algorithm 4. For a more detailed introduction into the VAE framework and its usage for CME parameterization, we refer the reader to [32].
(a) QPSK
(b) 16QAM
(c) Gaussian Symbols
Figure 3: NMSE over the SNR for given channel estimations in a $J=8$ user and $M=64$ antennas scenario based on $N=200$ observations, including $N_{p}=J$ pilots (if not stated otherwise) and (a) QPSK, (b) 16QAM, and (c) Gaussian data symbols, using measurement data (Sec. II-B).
IV-C Complexity Analysis
The standalone GMM estimator proposed by [30] precomputes the filters used for the individual components, resulting in a complexity of $\mathcal{O}(KM^{2})$. For the standalone VAE, the complexity is given as $\mathcal{O}(DM^{2})$ [32], where $D$ denotes the number of layers in the VAE’s forward pass. For our semi-blind methods, the subspace calculation requires $\mathcal{O}((N+J)M^{2})$. This results from calculating the sample covariance matrix with $\mathcal{O}(NM^{2})$ and taking the eigenvectors of the $J$ largest eigenvalues for the solution of (34). Using the projection approximation subspace tracking (PAST) algorithm [42], the computational complexity of the subspace calculation reduces to $\mathcal{O}(JM)$ for every update. In the case of the subspace GMM, the $K$ LMMSE estimates can not be precomputed, which results in a complexity of $\mathcal{O}(K(M^{2}+JM^{2}+J^{3}))$. Similarily, the subspace VAE exhibits a complexity of $\mathcal{O}(DM^{2}+JM^{2}+J^{3})$. For the projected versions of the GMM and VAE the complexity becomes $\mathcal{O}(KM^{2}+JM^{2})$ and $\mathcal{O}(DM^{2}+JM^{2})$, respectively. One should note that the calculation for each of the $K$ components in the GMM can be parallelized. Similarly, the computations in the VAE’s convolutional layers can be parallelized, mitigating the complexity.
V Baseline estimators
The following baseline channel estimators are considered for comparison with our methods. Based on the found subspace $\mathrm{range}({\bm{V}})$ we can formulate the pilot-based ML estimator as $\hat{{\bm{h}}}_{\text{ML}}={\bm{V}}{\bm{V}}^{\mathrm{H}}{\bm{y}}_{p}$, which is the closed-form solution to (16). This can be interpreted as the subspace-adjusted version of the conventional least squares (LS) channel estimator given as $\hat{{\bm{h}}}_{\text{LS}}={\bm{y}}_{p}$.
Another estimator is based on the sample covariance matrix, which we can compute from the training data set $\mathcal{H}$} to infer the channels’ global statistics as
${\bm{C}}_{s}=\frac{1}{|{\mathcal{H}}|}\sum_{{{\bm{h}}}\in{\mathcal{H}}}{{\bm{h}}}{{\bm{h}}}^{\mathrm{H}}.$ (60)
We can use the matrix ${\bm{C}}_{s}$ as a statistical prior to parameterize the semi-blind channel estimators outlined in Section III-A and Section III-B as
$\hat{{\bm{h}}}_{\text{sub. s-cov}}={\bm{V}}{\bm{V}}^{\mathrm{H}}{\bm{C}}_{s}{\bm{V}}\left({\bm{V}}^{\mathrm{H}}{\bm{C}}_{s}{\bm{V}}+\sigma^{2}\mathbf{I}_{J}\right)^{-1}{\bm{V}}^{\mathrm{H}}{\bm{y}}_{p},$ (61)
and
$\hat{{\bm{h}}}_{\text{proj. s-cov}}={\bm{C}}_{s}\left({\bm{C}}_{s}+{\bm{C}}_{\tilde{{\bm{n}}}}\right)^{-1}{\bm{P}}_{\bm{H}}{\bm{y}}_{p}.$ (62)
To evaluate the gain achieved by the semi-blind adaptions of the CGLM channel estimators, we also compare to the pilot-only GMM [30] and VAE [32].
Lastly, we compare our proposed methods to two iterative algorithms optimizing the ML formulation in (13), namely the EM from [15] and a MP variant similar to [16], which we run both until convergence or $500$ iterations, whichever comes first.
VI Numerical Simulations
(a) Spatial Channel Model (Sec. II-A)
(b) Measurement Data (Sec. II-B)
Figure 4: NMSE over the SNR for given channel estimations in a $J=8$ user and $M=64$ antennas scenario based on $N=200$ observations, including $N_{p}=J$ pilots and Gaussian symbols, using (a) the spatial channel model and (b) measurement data.
Figure 5: NMSE over the number of antennas for given channel estimations in a $J=M/8$ user scenario based on $N=25J$ observations, including $N_{p}=J$ pilots and Gaussian symbols at SNR =0=0 dB, using the spatial channel model.
To evaluate our proposed methods, we use channel realizations, which are normalized with $\mathbb{E}\left[\|{\bm{h}}\|^{2}\right]=M$. Thus, we can define the SNR as $\sigma^{-2}$. Further, the normalized MSE (MSE) defined as
$\text{NMSE}=\frac{1}{ML}\sum_{\ell=1}^{L}\|{\bm{h}}_{\ell}-\hat{{\bm{h}}}_{\ell}\|^{2},$ (63)
is used to characterize the estimators’ performances based on $L=10^{3}$ unseen channel samples stemming from the channel models detailed in Sections II-A and II-B. We use $\mathcal{H}=\{{\bm{h}}_{t}\}^{T}_{t=1}$ with $T=1.5\cdot 10^{5}$ training samples from the respective channel model to train the GMM and VAE, where we set the number of components to $K=64$ and the latent dimension to $Z=32$, respectively. Further, in the case of the VAE we allow non-zero values for $\bm{\mu}_{\bm{\theta}}({\bm{z}})$, as we use the circulant approximation for the spatial channel model and the block-Toeplitz property is not perfectly fulfilled for the measurement data, due to hardware imperfections. The VAE’s remaining implementation details are the same as provided in [32]. The assumption of spatial uncorrelated channels, as used in (23), only holds for the spatial channel model of Section II-A. For the case of the measurement campaign described in Section II-B, we approximate the noise covariance matrix in (22) as
${\bm{C}}_{\tilde{{\bm{n}}}}\approx\sigma^{2}\frac{J}{M}\mathbf{I}_{M}.$ (64)
Using the $T$ training samples to empirically approximate the projected noise covariance matrix (22) did not result in a different performance compared to using the approximation (64). The “s-cov” variants (“sub. s-cov” and “proj. s-cov”) utilize the same training samples. For most of the simulations, the number of BS antennas is set to $M=64$, cf. Sections II-A and II-B, serving $J=8=M/8$ number of users, a representative operating point [38, Chap. 1.3.3]. Further, if not stated otherwise, the number of snapshots is set to $N=200$, corresponding to a scenario that allows high channel dispersion and mobility, e.g., up to $135$ kph, c.f. [38, Chap. 2.1].
In Fig. 3, the channel estimation performances are compared for different symbol constellations using the proposed GMM based semi-blind channel estimation variants. Additionally, we compare to the purely pilot-based GMM channel estimator with $N_{p}=J$ and $N_{p}=N$ number of pilots, where the latter assumes all sent symbols to be known by the receiver. We utilize Gaussian symbols with $x_{j}(n)\sim\mathcal{N}_{\mathbb{C}}(0,P_{j}=1/J)$ such that $\sum_{j=1}^{J}P_{j}=1$ as well as QPSK and 1616QAM. In the latter two cases, also a decision directed GMM is evaluated, which decodes the sent symbols based on a pilot-based GMM channel estimate and uses the decoded symbols as additional pilots in a second stage. For the decoding, we utilize LMMSE equalization with a subsequent mapping to the closed constellation point.
Firstly, we see no qualitative difference between the different symbol constellations. Using a continuous symbol constellation has a negligible effect on the simulation results, as also previously observed in [15] and, hence, for the rest of this work, we utilize Gaussian symbols. Further, we see in Fig. 3 an improvement from the pilot-only GMM to our proposed semi-blind variants, showcasing again the advantage of utilizing the payload symbols in channel estimation. The full pilot GMM using $N_{p}=N$ pilot symbols achieves the SNR improvement as shown in Section II. The decision directed GMM based estimator saturates for high SNR values, making it an inferior semi-blind strategy.
(a) Spatial Channel Model (Sec. II-A)
(b) Measurement Data (Sec. II-B)
Figure 6: NMSE over the number of observations for given channel estimations in a $J=8$ user and $M=64$ antennas scenario based on NN observations, including $N_{p}=J$ pilots and Gaussian symbols at SNR =0=0 dB, using (a) the spatial channel model and (b) measurement data.
Fig. 4a and Fig. 4b show the different channel estimators’ performances with respect to the SNR for the spatial channel model (cf. Section II-A) and measurement data (cf. Section II-B), respectively. One can see that the semi-blind methods utilizing the CGLMs perform the best across the whole SNR. The projected variants of the CGLMs slightly outperform their subspace counterpart for most SNR values, which follows the derivations in Section III. Interestingly, the order of the semi-blind GMM and semi-blind VAE depends on the utilized channel model. In Fig. 4a, the projected GMM and VAE show both the best overall result, whereas in Fig. 4b, the projected GMM outperforms the VAE-based estimator. Additionally, we see that in Fig. 4a, the subspace VAE outperforms the subspace GMM, and in Fig. 4b, the results are vice versa. This ordering follows the ordering of the standalone version, where the plain GMM is better than the plain VAE in the case of the measurement data and worse for the spatial channel model. This comes from the fact that the Toeplitz assumption at the VAE may not be fulfilled due to impairments in the real measurement data. For high SNR values, all semi-blind variants approach each other except the EM and MP methods. In Fig. 4a, the EM and MP drastically improve from 1515dB to 2020dB, showing similar performance at 2020dB as the other semi-blind methods, whereas, in Fig. 4b, they show inferior results also for high SNR. The CGLM-based approaches keep a slight advantage even for high SNR values, which can be attributed to the fact that prior information is beneficial even for high SNR. A notable observation is that, in the mid-SNR range, the semi-blind CGLM variants outperform all related estimators by roughly $3$ dB.
In Fig. 5 we evaluate the respective estimators across different numbers of antennas at the BS. In order to make the comparison fair, we keep the number of users at a constant ratio of $J=M/8$ and set the number of snapshots to $N=25J$. With this, we only see an improvement for higher numbers of antennas for the CGLMs, while the other methods roughly keep their performance. Due to the constant parameter ratio, the VAE variants’ drastic performance gains for a higher number of antennas solely come from the VAE’s strong capability of modeling high-dimensional data.
[…]
A critical decrease in performance can be observed for the EM and MP algorithms. Here, the NMSE increases after a certain point when increasing the number of snapshots. Even though the minimum appears at different $N$, the overall behavior exhibits similarities. This is because both methods optimize the joint ML formulation in (13), where the optimization of the second term becomes dominant for a high number of snapshots. Hence, the pilot observation’s impact, which is relevant to estimating the channel phase, vanishes.
(a) Spatial Channel Model (Sec. II-A)
(b) Measurement Data (Sec. II-B)
Figure 7: NMSE over the number of users for given channel estimations in a $M=64$ antennas scenario based on $N=200$ observations, including $N_{p}=J$ pilots and Gaussian symbols at SNR =0=0 dB, using (a) the spatial channel model and (b) measurement data.
[…]
Overall, we can conclude that the proposed semi-blind CGLMs show superior channel estimation performance across all different setups. Depending on the used channel model, either the semi-blind GMMs or the semi-blind VAEs result in slightly better NMSE, where only in the case of the spatial channel model, the subspace GMM shows slightly worse performance compared to the other proposed methods. Moreover, the projected CGLMs outperform their respective subspace counterpart for most simulated operating points, showing the proposed projection method’s superiority.
VII Conclusion
This work presented a novel semi-blind channel estimation technique based on the class of CGLMs. To this end, two methods that incorporate subspace knowledge into the well-known LMMSE estimator are discussed. Both methods exploit the estimated subspace derived from the dominant eigenvectors of sample covariance matrices constructed using the received symbols. A theoretical analysis of the methods showed the proposed projection-based estimator’s superior estimation quality for uncorrelated Rayleigh fading channels. Secondly, we showed how two examples from the class of CGLMs, i.e., the GMM and VAE, can be used to parameterize these estimators. Extensive simulations based on real-world measurement and spatial channel model data demonstrated the proposed methods’ superior estimation performances compared to standard semi-blind channel estimators.
-A MSE of Projected LMMSE
For any linear estimator $\hat{{\bm{h}}}={\bm{W}}{\bm{y}}_{p}$, the MSE is given as
$\mathrm{MSE}$ $=\mathbb{E}\left[\|{\bm{h}}-\hat{{\bm{h}}}\|^{2}\right]=\mathbb{E}\left[\mathrm{tr}\left(\left({\bm{h}}-\hat{{\bm{h}}}\right)\left({\bm{h}}^{\mathrm{H}}-\hat{{\bm{h}}}^{\mathrm{H}}\right)\right)\right]$ (65)
$=\mathbb{E}\left[\mathrm{tr}({\bm{h}}{\bm{h}}^{\mathrm{H}})-2\mathrm{tr}({\bm{h}}{\bm{y}}^{\mathrm{H}}{\bm{W}}^{\mathrm{H}})+\mathrm{tr}({\bm{W}}{\bm{y}}{\bm{y}}^{\mathrm{H}}{\bm{W}}^{\mathrm{H}})\right].$ (66)
For the case of the projected LMMSE the second term in (66) can be rewritten as
$\mathbb{E}\left[\mathrm{tr}\left({\bm{h}}\tilde{{\bm{y}}}^{\mathrm{H}}{\bm{W}}^{\mathrm{H}}\right)\right]$ (67)
$\quad\quad=\mathbb{E}\left[\mathrm{tr}\left({\bm{h}}{\bm{h}}^{\mathrm{H}}{\bm{W}}^{\mathrm{H}}\right)\right]$ (68)
$\quad\quad=\mathbb{E}\left[\mathrm{tr}\left({\bm{h}}{\bm{h}}^{\mathrm{H}}\left({\bm{C}}+\sigma^{2}\frac{J}{M}\mathbf{I}_{M}\right)^{-1}{\bm{C}}\right)\right]$ (69)
$\quad\quad=\mathrm{tr}\left({\bm{C}}\left({\bm{C}}+\sigma^{2}\frac{J}{M}\mathbf{I}_{M}\right)^{-1}{\bm{C}}\right).$ (70)
Similarly for the third term in (66) we have
$\mathbb{E}\left[\mathrm{tr}\left({\bm{W}}\tilde{{\bm{y}}}\tilde{{\bm{y}}}^{\mathrm{H}}{\bm{W}}^{\mathrm{H}}\right)\right]$ (71)
$\quad\quad=\mathbb{E}\left[\mathrm{tr}\left({\bm{W}}\left({\bm{h}}{\bm{h}}^{\mathrm{H}}+\tilde{{\bm{n}}}\tilde{{\bm{n}}}^{\mathrm{H}}\right){\bm{W}}^{\mathrm{H}}\right)\right]$ (72)
$\quad\quad=\mathbb{E}\bigg[\mathrm{tr}\bigg({\bm{C}}\left({\bm{C}}+\sigma^{2}\frac{J}{M}\mathbf{I}_{M}\right)^{-1}\left({\bm{h}}{\bm{h}}^{\mathrm{H}}+\tilde{{\bm{n}}}\tilde{{\bm{n}}}^{\mathrm{H}}\right)$
$\quad\quad\quad\times\left({\bm{C}}+\sigma^{2}\frac{J}{M}\mathbf{I}_{M}\right)^{-1}{\bm{C}}\bigg)\bigg]$ (73)
$\quad\quad=\mathrm{tr}\left({\bm{C}}\left({\bm{C}}+\sigma^{2}\frac{J}{M}\mathbf{I}_{M}\right)^{-1}{\bm{C}}\right),$ (74)
where we assume that $\mathbb{E}\left[\tilde{{\bm{n}}}\tilde{{\bm{n}}}^{\mathrm{H}}\right]=\sigma^{2}\frac{J}{M}\mathbf{I}_{M}$. From this, the overall MSE in (26) follows directly.
-B MSE of Subspace LMMSE for Rayleigh Fading
In the case of uncorrelated Rayleigh fading, the subspace LMMSE filter is given as
${\bm{W}}_{\mathrm{sub}}=\frac{1}{1+\sigma^{2}}{\bm{V}}{\bm{V}}^{\mathrm{H}}.$ (75)
Using this filter the second term in (66) can be rewritten as
$\mathbb{E}\left[\mathrm{tr}\left({\bm{h}}{\bm{y}}^{\mathrm{H}}{\bm{W}}_{\mathrm{sub}}^{\mathrm{H}}\right)\right]$ (76)
$\quad\quad=\mathbb{E}_{\bm{h}}\left[\mathbb{E}\left[\mathrm{tr}\left({\bm{h}}{\bm{h}}^{\mathrm{H}}{\bm{W}}_{\mathrm{sub}}^{\mathrm{H}}\right)\mid{\bm{h}}\right]\right]$ (77)
$\quad\quad=\frac{1}{1+\sigma^{2}}\mathbb{E}_{\bm{h}}\left[\mathbb{E}\left[\mathrm{tr}\left({\bm{h}}{\bm{h}}^{\mathrm{H}}{\bm{V}}{\bm{V}}^{\mathrm{H}}\right)\mid{\bm{h}}\right]\right]$ (78)
$\quad\quad=\frac{1}{1+\sigma^{2}}\mathbb{E}_{\bm{h}}\left[\mathrm{tr}\left({\bm{h}}{\bm{h}}^{\mathrm{H}}\right)\right]$ (79)
$\quad\quad=\frac{1}{1+\sigma^{2}}M.$ (80)
Similarly for the third term in (66) we have
$\mathbb{E}\left[\mathrm{tr}\left({\bm{W}}_{\mathrm{sub}}{\bm{y}}{\bm{y}}^{\mathrm{H}}{\bm{W}}_{\mathrm{sub}}^{\mathrm{H}}\right)\right]$ (81)
$\quad\quad=\mathbb{E}_{\bm{h}}\left[\mathbb{E}\left[\mathrm{tr}\left({\bm{W}}_{\mathrm{sub}}{\bm{h}}{\bm{h}}^{\mathrm{H}}{\bm{W}}_{\mathrm{sub}}^{\mathrm{H}}\right)\mid{\bm{h}}\right]\right]$
$\quad\quad\quad+\mathbb{E}_{\bm{h}}\left[\mathbb{E}\left[\mathrm{tr}\left({\bm{W}}_{\mathrm{sub}}{\bm{n}}{\bm{n}}^{\mathrm{H}}{\bm{W}}_{\mathrm{sub}}^{\mathrm{H}}\right)\mid{\bm{h}}\right]\right]$ (82)
$\quad\quad=\frac{1}{(1+\sigma^{2})^{2}}\Big[\mathbb{E}_{\bm{h}}\left[\mathbb{E}\left[\mathrm{tr}\left({\bm{V}}{\bm{V}}^{\mathrm{H}}{\bm{h}}{\bm{h}}^{\mathrm{H}}{\bm{V}}{\bm{V}}^{\mathrm{H}}\right)\mid{\bm{h}}\right]\right]$
$\quad\quad\quad+\mathbb{E}_{\bm{h}}\left[\mathbb{E}\left[\mathrm{tr}\left({\bm{V}}{\bm{V}}^{\mathrm{H}}{\bm{n}}{\bm{n}}^{\mathrm{H}}{\bm{V}}{\bm{V}}^{\mathrm{H}}\right)\mid{\bm{h}}\right]\right]\Big]$ (83)
$\quad\quad=\frac{1}{(1+\sigma^{2})^{2}}\left[\mathbb{E}_{\bm{h}}\left[\mathrm{tr}\left({\bm{h}}{\bm{h}}^{\mathrm{H}}\right)\right]+\mathbb{E}_{\bm{h}}\left[\sigma^{2}\mathrm{tr}\left({\bm{V}}{\bm{V}}^{\mathrm{H}}\right)\right]\right]$ (84)
$\quad\quad=\frac{1}{(1+\sigma^{2})^{2}}(M+J\sigma^{2}).$ (85)
The overall subspace variant’s MSE for ${\bm{C}}=\mathbf{I}_{M}$ is then
$\mathrm{MSE}_{\mathrm{iid}}^{\mathrm{sub}}$ $=M-2\frac{1}{1+\sigma^{2}}M+\frac{1}{(1+\sigma^{2})^{2}}(M+J\sigma^{2})$ (86)
$=\frac{\sigma^{2}(M\sigma^{2}+J)}{(1+\sigma^{2})^{2}}.$ (87)
-C Derivation of Bayesian Cramer-Rao Bounds
The BCRB, which is the lower bound on the MSE of any estimator $\hat{{\bm{h}}}({\bm{Y}})$, is given as [48]
$\mathbb{E}\left[\|\hat{{\bm{h}}}({\bm{Y}})-{\bm{h}}\|^{2}\right]\geq\mathrm{tr}\left(\left[\left(\mathbb{E}_{{{\bm{h}}}}[\mathcal{I}({{{\bm{h}}}})]+\bm{J}_{p}\right)^{-1}\right]_{:M,:M}\right),$ (88)
where $\mathcal{I}({{{\bm{h}}}})$ is the Fisher information matrix corresponding to the inverse of the unbiased CRB, and
$\bm{J}_{p}$ $=\mathbb{E}_{\bm{h}}\left[\frac{\partial\log p({\bm{h}})}{\partial{\bm{h}}^{\mathrm{H}}}\frac{\partial\log p({\bm{h}})}{\partial{\bm{h}}}\right]$ (89)
$=\mathrm{blckdiag}({{\bm{C}}}_{1},\dots,{{\bm{C}}}_{J})^{-1},$ (90)
We denote with $\cdot]_{:M,:M}$ in (88) the part of the overall BCRB matrix corresponding to the upper left block of size $M\times M$, where we assume without loss of generality to index the users such that the user of interest is the first one.
For the case of ${\bm{C}}_{j}=\mathbf{I}_{M}$ for all $j\in\{1,\dots,J\}$ the prior information $\bm{J}_{p}$ is given as
$\bm{J}_{p}=\mathbf{I}_{JM}.$ (91)
For the deterministic BCRB, we use the formulation derived in [40], which results in
$\mathbf{BCRB}_{\mathrm{iid,d}}$
$\quad=\sigma^{2}\left(({\bm{X}}^{*}{\bm{X}}^{\mathrm{T}})\otimes{\bm{P}}_{\bm{H}}^{\perp}+({\bm{P}}^{*}{\bm{P}}^{\mathrm{T}})\otimes{\bm{P}}_{\bm{H}}+\sigma^{2}\mathbf{I}_{JM}\right)^{-1},$ (92)
with ${\bm{P}}_{\bm{H}}^{\perp}=\mathbf{I}_{M}-{\bm{P}}_{\bm{H}}$.
Now using the Fisher information matrix for the stochastic CRB as provided in [15], we introduce $\bar{{\bm{h}}}=[\Re\{{\bm{h}}_{1}^{\mathrm{T}}\},\Im\{{\bm{h}}_{1}^{\mathrm{T}}\},\dots,\Re\{{\bm{h}}_{J}^{\mathrm{T}}\},\Im\{{\bm{h}}_{J}^{\mathrm{T}}\}]^{\mathrm{T}}$, which gives us the prior information as
$\bm{J}_{p}$ $=\mathbb{E}_{\bar{{\bm{h}}}}\left[\frac{\partial\log p(\bar{{\bm{h}}})}{\partial\bar{{\bm{h}}}^{\mathrm{T}}}\frac{\partial\log p(\bar{{\bm{h}}})}{\partial\bar{{\bm{h}}}}\right]$ (93)
$=2\,\mathrm{blckdiag}(\bar{{\bm{C}}}_{1},\dots,\bar{{\bm{C}}}_{J})^{-1},$ (94)
with
$\bar{{\bm{C}}}_{j}=\begin{bmatrix}\Re\{{\bm{C}}_{j}\}&-\Im\{{\bm{C}}_{j}\}\\ \Im\{{\bm{C}}_{j}\}&\Re\{{\bm{C}}_{j}\}\end{bmatrix}$. (95)
Thus, the BCRB for uncorrelated Rayleigh fading is given as
$\mathbf{BCRB}_{\mathrm{iid,s}}=\left(\left(\frac{2N_{p}}{J\sigma^{2}}+2\right)\mathbf{I}_{2JM}+{\bm{R}}\right)^{-1},$ (96)
where $\bm{R}$} denotes the part depending on the payload symbols defined in [15].
-D Proof of Theorem 2
To this end, let us consider the MSE for a fixed channel matrix $\bm{H}$}, where the expectation is only taken over the noise and Gaussian symbol realizations as
$\mathbb{E}_{{\bm{X}},{\bm{N}}}[\|\hat{{\bm{h}}}_{\mathrm{proj}}(\hat{{\bm{V}}})-\hat{{\bm{h}}}_{\mathrm{proj}}({{\bm{V}}})\|^{2}]$
$\quad=\mathbb{E}_{{\bm{X}},{\bm{N}}}[\|{\bm{W}}_{\mathrm{proj}}(\hat{{\bm{V}}}\hat{{\bm{V}}}^{\mathrm{H}}-{\bm{V}}{\bm{V}}^{\mathrm{H}}){\bm{y}}\|^{2}]$ (97)
$\quad\leq\lambda_{\max}({\bm{W}}_{\mathrm{proj}}^{\mathrm{H}}{\bm{W}}_{\mathrm{proj}})\mathbb{E}_{{\bm{X}}}\left[(\mathrm{tr}({\bm{A}})+\mathbb{E}_{{\bm{N}}}[\mathrm{tr}({\bm{B}})])\right]$ (98)
with
$\mathrm{tr}({\bm{A}})$ $=\mathrm{tr}((\hat{{\bm{V}}}\hat{{\bm{V}}}^{\mathrm{H}}-{\bm{V}}{\bm{V}}^{\mathrm{H}}){\bm{h}}{\bm{h}}^{\mathrm{H}}(\hat{{\bm{V}}}\hat{{\bm{V}}}^{\mathrm{H}}-{\bm{V}}{\bm{V}}^{\mathrm{H}}))$ (99)
$=\mathrm{tr}((\mathbf{I}-\hat{{\bm{V}}}\hat{{\bm{V}}}^{\mathrm{H}}){\bm{h}}{\bm{h}}^{\mathrm{H}}(\mathbf{I}-\hat{{\bm{V}}}\hat{{\bm{V}}}^{\mathrm{H}}))$ (100)
$=\|(\mathbf{I}-\hat{{\bm{V}}}\hat{{\bm{V}}}^{\mathrm{H}}){\bm{h}}\|^{2}$ (101)
$=\|(\mathbf{I}-\hat{{\bm{V}}}\hat{{\bm{V}}}^{\mathrm{H}}){\bm{V}}{\bm{s}}\|^{2}$ (102)
$=\sum_{i\leq J}\sum_{j>J}|\hat{{\bm{v}}}_{j}^{\mathrm{H}}{\bm{v}}_{i}s_{i}|^{2}$ (103)
$\leq|s_{\max}|^{2}\sum_{j\leq J}\sum_{i>J}|\hat{{\bm{v}}}_{j}^{\mathrm{H}}{\bm{v}}_{i}|^{2}$ (104)
$=|s_{\max}|^{2}\sum_{i\leq J}\sum_{j>J}|\hat{{\bm{v}}}_{j}^{\mathrm{H}}{\bm{v}}_{i}|^{2}$ (105)
where ${\bm{h}}={\bm{V}}{\bm{s}}$ with ${\bm{s}}=[s_{1},\dots,s_{J}]^{\mathrm{T}}$ is the channel of interest’s decomposition in terms of $\bm{V}$}, and $s_{\max}=\max_{i}s_{i}$. Furthermore, we have
$\mathbb{E}_{\bm{N}}[\mathrm{tr}({\bm{B}})]$ $=\mathbb{E}_{\bm{N}}[\mathrm{tr}((\hat{{\bm{V}}}\hat{{\bm{V}}}^{\mathrm{H}}-{\bm{V}}{\bm{V}}^{\mathrm{H}}){\bm{n}}{\bm{n}}^{\mathrm{H}}(\hat{{\bm{V}}}\hat{{\bm{V}}}^{\mathrm{H}}-{\bm{V}}{\bm{V}}^{\mathrm{H}}))]$ (106)
$=\sigma^{2}(\mathrm{tr}(\hat{{\bm{V}}}\hat{{\bm{V}}}^{\mathrm{H}}+{\bm{V}}{\bm{V}}^{\mathrm{H}})-2\mathrm{tr}(\hat{{\bm{V}}}\hat{{\bm{V}}}^{\mathrm{H}}{\bm{V}}{\bm{V}}^{\mathrm{H}}))$ (107)
$=2\sigma^{2}\left[J-\sum_{i,j\leq J}|\hat{{\bm{v}}}_{j}^{\mathrm{H}}{\bm{v}}_{i}|^{2}\right]$ (108)
$=2\sigma^{2}\sum_{j\leq J}\sum_{i>J}|\hat{{\bm{v}}}_{j}^{\mathrm{H}}{\bm{v}}_{i}|^{2}$ (109)
Thus,
$\mathbb{E}_{{\bm{N}}}[\|\hat{{\bm{h}}}_{\mathrm{proj}}(\hat{{\bm{V}}})-\hat{{\bm{h}}}_{\mathrm{proj}}({{\bm{V}}})\|^{2}]$
$\quad\leq\lambda_{\max}({\bm{W}}_{\mathrm{proj}}^{\mathrm{H}}{\bm{W}}_{\mathrm{proj}})\left(|s_{\max}|^{2}+2\sigma^{2}\right)\sum_{j\leq J}\sum_{i>J}|\hat{{\bm{v}}}_{j}^{\mathrm{H}}{\bm{v}}_{i}|^{2}.$ (110)
Now, let us consider the following corollary from [49, Corollary 4.1] about the inner product of eigenvectors of sample and true covariance matrices.
Corollary 2.1.
Let us denote with $\lambda^{\prime}_{j}$ and ${\bm{v}}_{j}$ the $i$-th eigenvalue and eigenvector of the true covariance matrix, respectively. Then, for any weights $w_{ji}$, which are non-zero when $\lambda^{\prime}_{i}\neq\lambda^{\prime}_{j}$ and $\mathrm{sgn}(\lambda^{\prime}_{i}-\lambda^{\prime}_{j})2\lambda^{\prime}_{i}>\mathrm{sgn}(\lambda^{\prime}_{i}-\lambda^{\prime}_{j})(\lambda^{\prime}_{i}+\lambda^{\prime}_{j})$, and real $\varepsilon>0$
$P\left(\sum_{j\leq J}\sum_{i>J}w_{ji}|\hat{{\bm{v}}}_{j}^{\mathrm{H}}{\bm{v}}_{i}|^{2}\leq\varepsilon\right)$
$\quad\geq 1-\sum_{j\leq J}\sum_{i>J}\frac{4k^{2}_{i}w_{ji}}{N\varepsilon(\lambda^{\prime}_{j}-\lambda^{\prime}_{i})^{2}},$ (111)
where $k_{i}^{2}=\mathbb{E}[\|{\bm{y}}{\bm{y}}^{\mathrm{H}}{\bm{v}}_{i}\|^{2}]-\lambda_{i}^{\prime,2}$.
In our given problem, the true covariance matrix is given as ${\bm{C}}_{{\bm{y}}\mid{\bm{H}}}={\bm{H}}{\bm{H}}^{\mathrm{H}}/J+\sigma^{2}\mathbf{I}_{M}$. It is then trivial to see that $\lambda^{\prime}_{i}=\sigma^{2},\forall i>J$. Furthermore it is shown that [49, Corollary 4.3]
$k_{i}^{2}=\lambda^{\prime}_{i}\left(\lambda^{\prime}_{i}+\mathrm{tr}(C_{{\bm{y}}\mid{\bm{H}}})\right).$ (112)
Hence, we can conclude
$P\left(\mathbb{E}_{{\bm{X}},{\bm{N}}}[\|\hat{{\bm{h}}}_{\mathrm{proj}}(\hat{{\bm{V}}})-\hat{{\bm{h}}}_{\mathrm{proj}}({{\bm{V}}})\|^{2}]\leq\varepsilon\right)$
$\geq 1-\frac{4k^{2}(M-J)\lambda_{\max}({\bm{W}}_{\mathrm{proj}}^{\mathrm{H}}{\bm{W}}_{\mathrm{proj}})}{J^{2}N\varepsilon}\sum_{j\leq J}\frac{|s_{\max}|^{2}+2\sigma^{2}}{\lambda_{j}^{2}},$ (113)
with $\lambda_{j}=(\lambda^{\prime}_{j}-\sigma^{2})/J$ for all $j\in\{1,\dots,J\}$ being the eigenvalues of ${\bm{H}}{\bm{H}}^{\mathrm{H}}$ and
$k^{2}=\sigma^{2}\left(\sigma^{2}+\mathrm{tr}(C_{{\bm{y}}\mid{\bm{H}}})\right).$ (114)